OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). A: y = −one thirdx − 6 B: y = −one thirdx − 14 C: y = 3x − 6 D: y = 3x − 14

OpenStudy (anonymous):

At the moment i believe the answer is -1/3x - 6 but im not really sure tho

OpenStudy (akashdeepdeb):

Do it JUST lLIKE the previous one!! :)

OpenStudy (akashdeepdeb):

Check your answer again! :)

OpenStudy (anonymous):

im doing it right now

OpenStudy (anonymous):

i tried it again and i got the answer C: y = 3x − 6 is parallel @AkashdeepDeb

OpenStudy (akashdeepdeb):

The answer is D: Check previously what I had done! :)

OpenStudy (anonymous):

how did u get 14 tho? @AkashdeepDeb

OpenStudy (akashdeepdeb):

Do you know the point-slope form of the line?

OpenStudy (anonymous):

3x?

OpenStudy (akashdeepdeb):

Yes the slope is 3 Because if you write the equation which you gave in the question as y=mx+b You get m=3!! :D

OpenStudy (anonymous):

but i still dont understand how u got 14

OpenStudy (akashdeepdeb):

If 2 lines have the SAME SLOPE they are parallel or are the same. SLOPE GIVES YOU HOW STEEP IT IS. 

OpenStudy (anonymous):

L,M,N Are parallel i understand that part

OpenStudy (akashdeepdeb):

Now we need a parallel line right? So the slope will be the same as 3 And also it is asked that it has to pass through the point (4,-2) So we use the Point-slope form of a line. It is very easy to derive [I'll do that after some time] The formula is: y-y1 = m(x-x1) And we already have m = 3 and (x1,y1) = (4,-2) Substitute these points in the formula and you'll get the equation of the parallel line!! :D Understood this till here? :)

OpenStudy (anonymous):

lol some what

OpenStudy (akashdeepdeb):

Did you go through that website for sure?

OpenStudy (anonymous):